OpenStudy (firejay5):

35 less than seven of a certain number is fifty more than 10 times of that number. What is the number?

OpenStudy (firejay5):

@Loser66

OpenStudy (anonymous):

So we have that 7x-35=50+10x Solving: 7x=85+10x -3x=85 x=-85/3... I dunno Check

OpenStudy (dan815):